Car A of mass 3000 kg moving at 72km/hr, east tries to overtake the car B of mass 1250 kg moving at 60km/hr, east. Car A collides with car B. The cars stick together after the collision. Find their velocities after the collision. Note: The two cars are moving in the same direction.

1
Expert's answer
2022-03-27T18:20:52-0400

The law of conservation of momentum gives

"m_1v_1+m_2v_2=(m_1+m_2)u"

The final velocity

"u=\frac{m_1v_1+m_2v_2}{m_1+m_2}"

"u=\frac{3000*72+1250*60}{3000+1250}=68\:\rm km/hr"